find the total surface area and the volume of a closed cylindrical container with radius 5cm and a height of 34cm
Answer by lwsshak3(11628)

find the total surface area and the volume of a closed cylindrical container with radius 5cm and a height of 34cm
***
total surface area=area of 2 ends+area of side
Area of both ends=2*π*5^2=50π
Area of side=circumference*height=2π*5*34=340π
total surface area=50π+340π=390π≈1225.22 sq cm
volume=area of end*height=25π*34=850π≈2670.35 cubic cm